OpenStudy (anonymous):

if x=-3 whats is tthe value of |x-4|-x squared A -8 B-2 C 7 D 16

OpenStudy (anonymous):

im guessing 16 because 3 squared is 9 and when you square root 9 you get 3 and 4 squared is 16 and when you square root it you get 16

OpenStudy (cwrw238):

|3-4| = |-1| = 1

OpenStudy (cwrw238):

(-3)^2 = 9

OpenStudy (anonymous):

so wouldn't it be D 16? because -4*- = 16 @cwrw238

OpenStudy (cwrw238):

oops sorry my first post was wrong its |-3-4| = |-7| = 7

OpenStudy (cwrw238):

the modulus lines make everything inside positive

OpenStudy (cwrw238):

no not D

OpenStudy (cwrw238):

7 - 9

OpenStudy (anonymous):

but what about the negative x squared

OpenStudy (cwrw238):

- (-3)^2 = -9

OpenStudy (cwrw238):